# Dormant BH candidates from Gaia DR3 summary diagnostics

### v1

@@
Dormant BH candidates from Gaia DR3 summary diagnostics
@@
(Submitted 7 October 2025)
We present a rigorous identification of candidates for dormant black holes (BHs) and neutron stars (NSs) in binaries using summary statistics from Gaia Data Release 3 (DR3), rather than full orbital solutions. Although Gaia astrometric orbits have already revealed a small sample of compact object binaries, many systems remain undetected due to stringent quality cuts imposed on the published orbits. Using a forward-modelling framework that simulates Gaia observables, in particular the renormalised unit weight error (ruwe) and radial velocity (RV) scatter, we infer posterior distributions for companion mass and orbital period via Markov chain Monte Carlo (MCMC) sampling, marginalising over nuisance orbital parameters. We validate our approach by comparing the predicted masses and periods against full orbit solutions from DR3, and by successfully recovering known compact object binaries as promising candidates. The method is best suited for systems with red giant primaries, which have more reliable Gaia RV scatter and a light centroid more likely dominated by one component, compared to main-sequence stars. And they are less likely to be triples with short-period inner binaries, which produce confounding signatures. We apply the method to three million giants and identify 556 systems with best-fit companion masses $\gtrsim 3\,\mathrm{M}_{\odot}$. Recovery simulations suggest our selection method is substantially more sensitive than the DR3 non-single-star catalogue, particularly for binaries with periods below 1 year and above $\sim 6$ years. These candidates represent promising targets for spectroscopic follow-up and Gaia DR4 analysis to confirm the presence of compact objects and build a more complete census of the Galactic BH and NS population. Candidate main-sequence stars with massive companions face a larger set of confounding effects. Therefore, we present an analogous catalogue of 279 additional ‘main sequence’ candidates only as an appendix.
@@ 1 Introduction
Stellar-mass black holes (BHs) are compact remnants of massive stars, formed through the gravitational collapse of their cores at the end of stellar evolution. Their masses, orbits, and binary companions encode crucial information about both their formation pathways, from the final stages of massive star evolution to supernova mechanisms and natal kicks. Their subsequent evolution may make them X-ray binaries or gravitational wave sources. Understanding the population demographics of stellar-mass BHs is thus essential for benchmarking theoretical models. However, our observational picture remains largely incomplete, with poorly constrained empirical connections between stellar progenitors (in terms of mass, metallicity, and binarity) and the resulting BH properties.
@@
Until recently, BHs were only detected in two sub-populations: X-ray binaries and gravitational wave events. Accreting BHs in close binaries can be identified through their high-energy X-ray emission, with approximately two dozen dynamically confirmed Galactic systems to date (e.g. Remillard & McClintock, 2006; Casares & Jonker, 2014; Corral-Santana et al., 2016). Merging BHs in double compact systems are detected via gravitational wave emission, with more than 200 candidate events reported in the latest LIGO/VIRGO/KAGRA observing run (LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA Collaboration et al., 2023). However, both these detection methods trace rare evolutionary pathways and hence sample only a tiny fraction of the total BH population ($\ll 1\%$, e.g. Wiktorowicz et al., 2019). Gravitational wave sources are observed only at cosmological distances and form through specific binary evolution scenarios, while only a small subset of BHs in binaries are actively accreting and luminous in X-rays. As a result, both methods provide a highly biased and incomplete view of the overall BH population.
@@
A much larger fraction of BHs is expected to exist either in wide binaries with luminous companions or as isolated objects. While gravitational microlensing provides a potential pathway for identifying isolated BHs, it requires rare alignments and has yielded only one good candidate to date (e.g. Lam et al., 2022; Sahu et al., 2022). Non-accreting ‘dormant’ BHs in binaries with stellar companions thus represent a particularly promising alternative avenue to expand the census of Galactic BHs. These systems are expected to be substantially more numerous than their accreting or merging counterparts, and can be detected through astrometric, kinematic, or photometric signatures induced by the BH on the companion (e.g. Shahaf et al., 2023; Jayasinghe et al., 2023; Gomel et al., 2023). The presence of a luminous companion further offers a unique opportunity to probe the properties and formation history of the BH through the observable characteristics of the surviving star.
@@
The third data release (DR3; Gaia Collaboration et al. 2023) of the Gaia mission (Gaia Collaboration et al., 2016) has opened up new possibilities for identifying dormant BHs in binaries (e.g. El-Badry, 2024, for a review on the Gaia binary renaissance). DR3 includes an extensive non-single-star (NSS) catalogue of binary orbits, derived from astrometric and spectroscopic measurements, with 169,227 and 220,372 published orbit solutions, respectively (Halbwachs et al., 2023; Gosset et al., 2025). This data set has already enabled the discovery of the first Galactic dormant BHs (El-Badry et al., 2023a; Chakrabarti et al., 2023; El-Badry et al., 2023b; Gaia Collaboration et al., 2024; Wang et al., 2024), as well as a dozen candidate neutron star (NS) companions (El-Badry et al., 2024b, c), all vetted by spectroscopic follow-up. These systems feature compact objects in wide orbits (typically $P\gtrsim 1$ yr) around low-mass stars. Their discovery underscores the power of Gaia to reveal a previously hidden population of compact remnants in the solar neighbourhood.
@@
However, the NSS catalogues with orbit solutions in Gaia DR3 were subject to stringent quality cuts (e.g. Halbwachs et al., 2023). For example, only astrometric binary solutions exceeding a strict signal-to-noise (S/N) threshold in parallax were included (Fig. 1). Notably, even the recently identified Gaia BH3 (Gaia Collaboration et al., 2024) – a $33\,\mathrm{M}_{\odot}$ BH orbiting a giant star– did not meet the S/N criteria for inclusion in the astrometric binary catalogue. Similarly, the candidate mass-gap BH Gaia DR3 3425577610762832384 (Wang et al., 2024) did not receive an astrometric solution. These cases highlight that many more BH and NS binaries may exist within close proximity to the Sun but remain unidentified due to the quality cuts for orbit solutions applied in DR3.
@@
Figure 1: The strict quality cuts for Gaia DR3 excised many binaries (some with BH companions) from the published Gaia NSS catalogue. The figure illustrates the cut on period vs. parallax over parallax uncertainty for binaries in the DR3 astrometric catalogue and the dormant BH binaries Gaia BH 1, 2 (Chakrabarti et al., 2023; El-Badry et al., 2023a, b).
@@
In addition to orbital solutions, Gaia DR3 provides time-averaged parameters and variability metrics for a much larger sample of stars. These statistics include the radial velocity (RV) uncertainty $\sigma_{\mathrm{RV}}$ (Katz et al., 2023), the renormalised unit weight error ruwe from the astrometric solution (Lindegren et al., 2018), and the photometric variability in the $G$, $G_{\mathrm{BP}}$, and $G_{\mathrm{RP}}$ bands $\sigma_{\mathrm{phot}}$ (Riello et al., 2021). The light centroid motion (in position or in RV) of unresolved binaries can produce excess scatter in these observables, providing indirect signatures of binarity even in the absence of a published orbit. These metrics are typically derived from 20–80 observational epochs and are available for 33 million stars (Katz et al., 2023).
@@
The light centroid’s orbital motion affects the stellar position on the sky, degrading the fit of the single-star astrometric model and leading to elevated ruwe values and astrometric excess noise. Spectroscopic binaries may exhibit inflated RV uncertainties and larger RV amplitudes, while close systems may show enhanced photometric variability due to ellipsoidal modulation or eclipses. These correlations have been widely exploited to identify unresolved binaries. In particular, ruwe has been shown to trace binarity across various regions of the colour-magnitude diagram (CMD; Belokurov et al., 2020; Penoyre et al., 2022b; Korol et al., 2022), and analytical models have been developed to predict its behaviour in binaries (e.g. Penoyre et al., 2022a; Andrew et al., 2022; Castro-Ginard et al., 2024). For systems with orbital periods shorter than the Gaia mission baseline, the expected ruwe signal approximately scales with the angular size of the photocentre orbit (Stassun & Torres, 2021), making ruwe most sensitive to nearby systems with intermediate periods ($P\sim 10^{3}$ days) and massive, faint companions.
@@
While each summary statistic on its own provides only weak constraints and may be influenced by other astrophysical variability (e.g. pulsations, spots, or multiple luminous stars), their joint modelling increases diagnostic power. Previous work by Andrew et al. (2022) combined astrometric, photometric, and spectroscopic variability indicators to search for compact object companions. Using analytic estimates of excess noise, they identified promising candidates in pre-DR3 data. A key challenge in their analysis was contamination from hierarchical triples, where large RV amplitudes and astrometric motion can arise from distinct companions in inner and outer orbits.
@@
Building on these efforts, we leverage the expanded dataset of Gaia DR3 and adopt a forward model of the full Gaia observing pipeline (El-Badry et al., 2024a). This approach allows us to interpret the observed combination of astrometric, spectroscopic, and photometric variability in a consistent and quantitative framework. Our model fits these observables simultaneously to constrain the mass $M_{\mathrm{CO}}$ and orbital period $P$ of possible binary companions. By selecting systems with high inferred companion masses and mass ratios, we both recover known compact object binaries and identify new candidates that may host dormant BHs or NSs. By focussing on evolved stars, where present or past large radii tend to exclude tight inner binaries, we reduce contamination from unresolved higher-order multiples.
@@
Sect. 2 gives an overview of the forward model framework we use to model Gaia summary diagnostics. In Sect. 3, we describe our approach for constraining binary parameters using these observables. The application of this method to evolved stars in Gaia DR3 is presented in Sect. 4, along with the resulting catalogue of candidate binaries with compact companions. We discuss potential sources of contamination in Sect. 5, and summarise our findings in Sect. 6.
2 Forward-modelling Gaia summary statistics
@@
In this section, we describe our forward modelling framework, which predicts the expected summary statistics reported by Gaia as a function of binary parameters. This modelling step is then used to solve the inverse problem: given a set of observed summary statistics, what physical parameters and orbits can explain them?
Figure 2: Predicted vs. catalogued Gaia DR3 summary statistics for known binaries, with the predictions based on forward modelling of their orbital parameters. Top panel: Comparison between predicted and catalogue values of the astrometric goodness of fit of the single-star astrometric solution, ruwe, for binaries with published orbital solutions. Bottom panel: Predicted vs. observed RV “errors” for spectroscopic binaries (SB1) in Gaia DR3. In both plots, the blue line shows the median trend, while the dashed blue lines indicate the 16th and 84th percentiles. The 1:1 relation is marked in black. Lower subpanels show the logarithmic residuals between observed and predicted values, with side histograms illustrating their normalised distribution. The mean bias ($\mu$) and scatter ($\sigma$) of the residuals are indicated.
@@ 2.1 Astrometry
We use the generative forward model gaiamock (El-Badry et al., 2024a) to simulate Gaia astrometric observations for unresolved binaries and predict the resulting summary statistics, particularly the renormalised unit weight error ruwe. gaiamock generates realistic epoch astrometry based on the Gaia scanning law, and the source parameters, including position ($\alpha,\delta$), parallax $\varpi$, proper motions ($\mu_{\alpha},\mu_{\delta}$), $G$ band magnitude, flux ratio, and component masses ($M_{*},M_{\mathrm{CO}}$). Alternatively, gaiamock can take as input the semi-major axis of the photocenter $a_{0}$ and the binary orbital elements: period $P$, time of periastron passage $T_{\mathrm{peri}}$, eccentricity $e$, inclination $i$, argument of periastron $\omega$, and longitude of ascending node $\Omega$. The model uses an empirical prescription for the epoch astrometry errors, calibrated on well-behaved single stars.
From the simulated epoch astrometry, gaiamock fits a five-parameter single-star model, predicts the resulting along-scan displacements, and computes the corresponding chi-squared ($\chi^{2}$) statistic of the residuals. Following Gaia conventions, the per-epoch uncertainties are inflated based on the goodness of fit, and ruwe is approximately the square root of this reduced chi-squared:
@@
Here, $\nu=\mathtt{astrometric\_n\_good\_obs\_al}-m$ is the number of degrees of freedom, calculated as the difference between the number of observations in the along-scan direction and the number of free parameters; $m=5$ for the five-parameter single-star model. To validate our astrometric model, we compared predicted ruwe values with those reported in Gaia DR3 for sources with published astrometric orbits (Halbwachs et al., 2023; Gaia Collaboration et al., 2023). These systems are well suited for validation, as their orbital parameters are fully constrained, enabling accurate forward modelling of their astrometric motion (this is in contrast to systems with only spectroscopic orbits, where, for example, orbital orientation is unconstrained, which would dominate the scatter in predicted ruwe).
We use binaries from the two_body_orbit catalogue with either astrometric-only (Orbital) or combined astrometric and spectroscopic (AstroSpectroSB1) solutions, selecting 158,690 high-quality orbits with $\texttt{goodness\_of\_fit}<10$ for $G\leq 13$ and <5<5 for $G>13$ (see El-Badry et al., 2023a). Fig. 2 shows that the predicted and catalogued ruwe values exhibit a tight correlation with negligible systematic offset, consistent with previous results (El-Badry, 2025). The scatter in the logarithmic ratio $\ln\left(\texttt{ruwe}_{\text{pred}}/\texttt{ruwe}_{\text{obs}}\right)$ is approximately constant $\epsilon_{\texttt{ruwe}}=0.14$ throughout the ruwe range. We adopt this value as the fractional uncertainty in our ruwe predictions.

### v2

# Dormant black hole candidates from Gaia DR3 summary diagnostics

@@
Dormant black hole candidates from Gaia DR3 summary diagnostics
@@
(Submitted 7 October 2025 / Accepted 26 February 2026)
We present a rigorous identification of candidates for dormant black holes (BHs) and neutron stars (NSs) in binaries using summary statistics from Gaia Data Release 3 (DR3), rather than full orbital solutions. Although Gaia astrometric orbits have already revealed a small sample of compact object binaries, many systems remain undetected due to stringent quality cuts imposed on the published orbits. Using a forward-modelling framework that simulates Gaia observables, in particular the re-normalised unit weight error (ruwe) and radial velocity (RV) scatter, we infer posterior distributions for companion mass and orbital period via Markov chain Monte Carlo (MCMC) sampling, marginalising over nuisance orbital parameters. We validate our approach by comparing the predicted masses and periods against full orbit solutions from DR3, and by successfully recovering known compact object binaries as promising candidates. The method is best suited for systems with red giant primaries, which have more reliable Gaia RV scatter and a light centroid more likely dominated by one component, compared to main-sequence stars, and they are less likely to be triples with short-period inner binaries, which produce confounding signatures. We applied the method to three million giants and identify 389 systems with best-fit companion masses $\gtrsim 3\,\mathrm{M}_{\odot}$. Recovery simulations suggest our selection method is substantially more sensitive than the DR3 non-single-star catalogue, particularly for binaries with periods below 1 year and above $\sim 6$ years. These candidates represent promising targets for spectroscopic follow-up and Gaia DR4 analysis to confirm the presence of compact objects and build a more complete census of the Galactic BH and NS population. Candidate main-sequence stars with massive companions face a larger set of confounding effects. Therefore, we present an analogous catalogue of 279 additional main-sequence candidates only as an appendix.
@@ 1 Introduction
Stellar-mass black holes (BHs) are compact remnants of massive stars, formed through the gravitational collapse of their cores at the end of stellar evolution. Their masses, orbits, and binary companions encode crucial information about their formation pathways, from the final stages of massive star evolution to supernova mechanisms and natal kicks. Their subsequent evolution may make them X-ray binaries or gravitational wave sources. Understanding the population demographics of stellar-mass BHs is thus essential for benchmarking theoretical models. However, our observational picture remains largely incomplete, with poorly constrained empirical connections between stellar progenitors (in terms of mass, metallicity, and binarity) and the resulting BH properties.
@@
Until recently, BHs had only been detected in two subpopulations: X-ray binaries and gravitational wave events. Accreting BHs in close binaries can be identified through their high-energy X-ray emission, and there have been approximately two dozen dynamically confirmed Galactic systems to date (e.g. Remillard and McClintock, 2006; Casares and Jonker, 2014; Corral-Santana et al., 2016). Merging BHs in double compact systems are detected via gravitational wave emission; more than 200 candidate events are reported in the latest LIGO/VIRGO/KAGRA observing run (LIGO Scientific Collaboration, Virgo Collaboration, and KAGRA Collaboration et al., 2023). However, both of these detection methods trace rare evolutionary pathways, and therefore sample only a tiny fraction of the total BH population ($\ll 1\%$; e.g. Wiktorowicz et al., 2019). Gravitational wave sources are observed only at cosmological distances and form through specific binary evolution scenarios, while only a small subset of BHs in binaries are actively accreting and are luminous in X-rays. As a result, both methods provide a highly biased and incomplete view of the overall BH population.
@@
A much larger fraction of BHs is expected to exist in wide binaries with luminous companions or as isolated objects. While gravitational microlensing provides a potential pathway for identifying isolated BHs, it requires rare alignments and has yielded very few good candidates to date (e.g. Lam et al., 2022; Sahu et al., 2022; Kruszyńska et al., 2024; Howil et al., 2025). Non-accreting dormant BHs in binaries with stellar companions thus represent a particularly promising alternative avenue to expand the census of Galactic BHs. These systems are expected to be substantially more numerous than their accreting or merging counterparts, and can be detected through astrometric, kinematic, or photometric signatures induced by the BH on the companion (e.g. Shahaf et al., 2023; Jayasinghe et al., 2023; Gomel et al., 2023). The presence of a luminous companion further offers a unique opportunity to probe the properties and formation history of the BH through the observable characteristics of the surviving star.
@@
The third data release (DR3; Gaia Collaboration et al. 2023b) of the Gaia mission (Gaia Collaboration et al., 2016) has opened up new possibilities to identify dormant BHs in binaries (e.g. El-Badry, 2024, for a review on the Gaia binary renaissance). DR3 includes an extensive non-single-star (NSS) catalogue of binary orbits, derived from astrometric and spectroscopic measurements, with 169,227 and 220,372 published orbit solutions, respectively (Halbwachs et al., 2023; Gosset et al., 2025). This dataset has already enabled the discovery of the first Galactic dormant BHs (El-Badry et al., 2023a; Chakrabarti et al., 2023; El-Badry et al., 2023b; Gaia Collaboration et al., 2024; Wang et al., 2024), as well as a dozen candidate neutron star (NS) companions (El-Badry et al., 2024b, c), all vetted by spectroscopic follow-up. These systems feature compact objects in wide orbits (typically $P\gtrsim 1$ yr) around low-mass stars. Their discovery underscores the power of Gaia to reveal a previously hidden population of compact remnants in the solar neighbourhood.
@@
However, the NSS catalogues with orbit solutions in Gaia DR3 were subject to stringent quality cuts (e.g. Halbwachs et al., 2023). For example, only astrometric binary solutions exceeding a strict signal-to-noise (S/N) threshold in parallax were included (Fig. 1). Notably, even the recently identified Gaia BH3 (Gaia Collaboration et al., 2024), a $33\,\mathrm{M}_{\odot}$ BH orbiting a giant star, did not meet the S/N criteria for inclusion in the astrometric binary catalogue. Similarly, the candidate mass-gap BH Gaia DR3 3425577610762832384 (Wang et al., 2024) did not receive an astrometric solution. These cases highlight that many more BH and NS binaries may exist within close proximity to the Sun, but remain unidentified due to the quality cuts for orbit solutions applied in DR3.
@@
Figure 1: Cut on period vs parallax over parallax uncertainty for binaries in the DR3 astrometric catalogue and the dormant BH binaries Gaia BH 1, 2 (Chakrabarti et al., 2023; El-Badry et al., 2023a, b). The strict quality cuts for Gaia DR3 excised many binaries (some with BH companions) from the published Gaia NSS catalogue.
@@
In addition to orbital solutions, Gaia DR3 provides time-averaged parameters and variability metrics for a much larger sample of stars. These statistics include the radial velocity (RV) uncertainty $\sigma_{\mathrm{RV}}$ (Katz et al., 2023), the re-normalised unit weight error ruwe from the astrometric solution (Lindegren, 2018), and the photometric variability in the $G$, $G_{\mathrm{BP}}$, and $G_{\mathrm{RP}}$ bands $\sigma_{\mathrm{phot}}$ (Riello et al., 2021). The light centroid motion (in position or in RV) of unresolved binaries can produce excess scatter in these observables, providing indirect signatures of binarity even in the absence of a published orbit. These metrics are typically derived from 20–80 observational epochs and are available for 33 million stars (Katz et al., 2023).
@@
The light centroid’s orbital motion affects the stellar position on the sky, degrading the fit of the single-star astrometric model and leading to elevated ruwe values and astrometric excess noise. Spectroscopic binaries may exhibit inflated RV uncertainties and larger RV amplitudes, while close systems may show enhanced photometric variability due to ellipsoidal modulation or eclipses. These correlations have been widely exploited to identify unresolved binaries. In particular, ruwe has been shown to trace binarity across various regions of the colour-magnitude diagram (CMD; Belokurov et al., 2020; Penoyre et al., 2022b; Korol et al., 2022), and analytical models have been developed to predict its behaviour in binaries (e.g. Penoyre et al., 2022a; Andrew et al., 2022; Castro-Ginard et al., 2024). For systems with orbital periods shorter than the Gaia mission baseline, the expected ruwe signal approximately scales with the angular size of the photocentre orbit (Stassun and Torres, 2021), making ruwe most sensitive to nearby systems with intermediate periods ($P\sim 10^{3}$ days) and massive faint companions.
@@
While each summary statistic on its own provides only weak constraints and may be influenced by other astrophysical variability (e.g. pulsations, spots, or multiple luminous stars), their joint modelling increases diagnostic power. Previous work by Andrew et al. (2022) combined astrometric, photometric, and spectroscopic variability indicators to search for compact object companions. Using analytic estimates of excess noise, they identified promising candidates in DR3 data. A key challenge in their analysis was contamination from hierarchical triples, where large RV amplitudes and astrometric motion can arise from distinct companions in inner and outer orbits.
@@
Building on these efforts, we leveraged the Gaia DR3 dataset and adopted a forward model of the entire Gaia observing pipeline (El-Badry et al., 2024a). This approach allowed us to interpret the observed combination of astrometric, spectroscopic, and photometric variability in a consistent and quantitative framework. Our model fits these observables simultaneously to constrain the mass $M_{\mathrm{CO}}$ and orbital period $P$ of possible binary companions. By selecting systems with high inferred companion masses and mass ratios, we recover known compact object binaries and identify new candidates that may host dormant BHs or NSs. By focussing on evolved stars, where present or past large radii tend to exclude tight inner binaries, we reduced the contamination from unresolved higher-order multiples.
@@
Section 2 gives an overview of the forward model framework we use to model Gaia summary diagnostics. In Sect. 3 we describe our approach to constraining binary parameters using these observables. The application of this method to evolved stars in Gaia DR3 is presented in Sect. 4, along with the resulting catalogue of candidate binaries with compact companions. We discuss potential sources of contamination in Sect. 5, and summarise our findings in Sect. 6.
2 Forward modelling Gaia summary statistics
@@
In this section we describe our forward-modelling framework, which predicts the expected summary statistics reported by Gaia as a function of binary parameters. This modelling step is then used to solve the inverse problem: given a set of observed summary statistics, what physical parameters and orbits can explain them?
Figure 2: Predicted vs catalogued Gaia DR3 summary statistics for known binaries. The predictions are based on forward modelling of their orbital parameters. Top panel: Comparison between predicted and catalogue values of the astrometric goodness of fit of the single-star astrometric solution, ruwe, for binaries with published orbital solutions. Bottom panel: Predicted vs observed RV errors for spectroscopic binaries (SB1) in Gaia DR3. In both plots the blue line shows the median trend, while the dashed blue lines indicate the 16th and 84th percentiles. The 1:1 relation is marked in black. The lower subpanels show the logarithmic residuals between observed and predicted values; the side histograms illustrate their normalised distribution. The mean bias ($\mu$) and scatter ($\sigma$) of the residuals are indicated.
@@ 2.1 Astrometry
We used the generative forward model gaiamock (El-Badry et al., 2024a) to simulate Gaia astrometric observations for unresolved binaries and predict the resulting summary statistics, particularly the re-normalised unit weight error ruwe. gaiamock generates realistic epoch astrometry based on the Gaia scanning law, and the source parameters, including position ($\alpha,\delta$), parallax $\varpi$, proper motions ($\mu_{\alpha},\mu_{\delta}$), $G$ band magnitude, flux ratio, and component masses ($M_{*},M_{\mathrm{CO}}$). Alternatively, gaiamock can take as input the semi-major axis of the photocenter $a_{0}$ and the binary orbital elements: period $P$, time of periastron passage $T_{\mathrm{peri}}$, eccentricity $e$, inclination $i$, argument of periastron $\omega$, and longitude of ascending node $\Omega$. The model uses an empirical prescription for the epoch astrometry errors, calibrated on well-behaved single stars.
From the simulated epoch astrometry, gaiamock fits a five-parameter single-star model, predicts the resulting along-scan displacements, and computes the corresponding chi-squared ($\chi^{2}$) statistic of the residuals. While a minority of Gaia DR3 sources are fitted with a six-parameter astrometric model (Lindegren et al., 2021b), these constitute a small fraction of our sample and were treated identically in the forward modelling with gaiamock. Following Gaia conventions, the per-epoch uncertainties are inflated based on the goodness of fit, and ruwe is approximately the square root of this reduced chi-squared (El-Badry, 2025):
@@
Here $\nu=\mathtt{astrometric\_n\_good\_obs\_al}-m$ is the number of degrees of freedom, defined as the difference between the number of observations in the along-scan direction and the number of free parameters; $m=5$ for the five-parameter single-star model. To validate our astrometric model, we compared predicted ruwe values with those reported in Gaia DR3 for sources with published astrometric orbits (Halbwachs et al., 2023; Gaia Collaboration et al., 2023b). These systems are well suited for validation, as their orbital parameters are fully constrained, enabling accurate forward modelling of their astrometric motion (this is in contrast to systems with only spectroscopic orbits, where, for example, orbital orientation is unconstrained, which would dominate the scatter in predicted ruwe).
We used binaries from the two_body_orbit catalogue with either astrometric-only (Orbital) or combined astrometric and spectroscopic (AstroSpectroSB1) solutions and selected 158,690 high-quality orbits with $\texttt{goodness\_of\_fit}<10$ for $G\leq 13$ and <5<5 for $G>13$ (see El-Badry et al., 2023a). Figure 2 shows that the predicted and catalogued ruwe values exhibit a tight correlation with negligible systematic offset, consistent with previous results (El-Badry, 2025). A minute fraction of outliers show predicted ruwe values larger than observed; these do not exhibit a clear correlation with any single orbital parameter. The scatter in the logarithmic ratio $\ln\left(\texttt{ruwe}_{\text{pred}}/\texttt{ruwe}_{\text{obs}}\right)$ is approximately constant $\epsilon_{\texttt{ruwe}}=0.14$ throughout the ruwe range. We adopt this value as the fractional uncertainty in our ruwe predictions.